Markus Company’s common stock sold for $1.50 per share at the end of this year. The company paid a common stock dividend of $0.39 per share this year. It also provided the following data excerpts from this year’s financial statements:     

Ending
Balance
Beginning
Balance
Cash $ 24,500 $ 40,800
Accounts receivable $ 42,000 $ 38,800
Inventory $ 42,400 $ 42,000
Current assets $ 108,900 $ 121,600
Total assets $ 302,000 $ 242,800
Current liabilities $ 46,500 $ 31,500
Total liabilities $ 92,000 $ 82,800
Common stock, $1 par value $ 107,000 $ 107,000
Total stockholders’ equity $ 210,000 $ 160,000
Total liabilities and stockholders’ equity $ 302,000 $ 242,800

     

This Year
Sales (all on account) $ 480,000
Cost of goods sold $ 278,400
Gross margin $ 201,600
Net operating income $ 32,500
Interest expense $ 2,000
Net income $ 21,350

1-1. What is the acid-test ratio at the end of this year?

1-2. What is the accounts receivable turnover and the average collection period?

1-3. What is the inventory turnover and the average sale period?

1-4. What is the company’s operating cycle?

1-5. What is the total asset turnover?

1-6. What is the times interest earned ratio?

2. What is the debt-to-equity ratio at the end of this year?

3. What is the equity multiplier?
Answer #1

1-1. Acid test ratio = Current assets - inventories / current liabilities

Acid test ratio = $108,900-42,400 / 46,500 = 1.43

1-2. Accounts receivable turnover = Credit sales / Average accounts receivable

Average accounts receivable = $38,800+42,000 / 2 = $40,400

Accounts receivable turnover = $480,000 / 40,400 = 11.88

Average collection period = 365 days / Accounts receivable turnover

Average collection period = 365 days / 11.88 = 30.72 days

1-3. Inventory turnover ratio = Cost of goods sold / Average inventory

Average inventory = $42,000+42,400 / 2 = $42,200

Inventory turnover ratio = $278,400 / 42,200 = 6.597

Average sale period = 365 days / Inventory turnover ratio

Average sale period = 365 days / 6.597 = 55.33 days

1-4. Operating cycle = 55.33 days + 30.72 days = 86.05 days

1-5. Total asset turnover = Sales / Average total asset

Average total asset = $242,800+302,000 / 2 = $272,400

Total asset turnover = $480,000 / 272,400 = 1.76

1-6. Times interest earned ratio = Earnings before interest and tax / interest expense

Times interest earned ratio = $32,500 / 2,000 = 16.25

2. Debt to equity ratio = Total liabilities / Total stockholders equity

Debt to equity ratio = $92,000 / 210,000 = 0.44

3. Equity multiplier = Average total assets / Average total stockholders equity

Average total stockholders equity = $160,000+210,000 / 2 = $185,000

Equity multiplier = $272,400 / 185,000 = 1.47